# Two-sample comparison through additive tree models for density ratios

### v2

@@ 2.3 Additive tree models and boosting
Given the close relationship between density ratios and posterior odds, the latter of which can be estimated using supervised tree boosting with additive tree models, it is natural to consider the class of additive tree models as a promising candidate for approximating the density ratio. We thus consider from now on fitting additive tree models to the density ratio under the balancing loss.
We begin with a non-Bayesian approach that aims to find a (functional) point estimate of the density ratio by optimizing the finite-sample version of the balancing loss. The resulting training algorithms are very similar to those used in supervised tree boosting. Before introducing the algorithms, we describe the additive tree model. We consider a model for the log-balancing weight $\log w$ in the form
@@
Proposition 2.1.
Suppose $w_{k-1}$ is the estimate of $w$ after (k1)(k-1) steps of fitting. Then in the $k$th step, the loss is minimized by $\log w_{k}=\log w_{k-1}+f_{k}$ when the following conditions are satisfied by $f_{k}$:
@@
    $e^{\beta_{k}(A)}=\sqrt{\frac{\int_{A}pw^{-1}_{k-1}d\mu}{\int_{A}qw_{k-1}d\mu}}.$
Remark 1: If the balancing weight is correctly estimated, it achieves the symmetry between the two expectations in the loss in that $\mathbb{E}_{p}[w^{-1}]=\mathbb{E}_{q}[w]$, and we can impose this condition as a restriction or regularization on the estimator so that it does not focus on minimizing only one expectation out of the two. As justification, Proposition 2.1 implies that the balancing loss can be improved by multiplying the current estimate $w_{k}$ by a constant ($c$, say) such that $\mathbb{E}_{p}[(cw_{k})^{-1}]=\mathbb{E}_{q}[cw_{k}]$ (see Supplementary Materials A for details). Based on this discussion, we introduce this correction step in the boosting algorithms.
@@
As in tree-based predictive models, searching for the exact optimal tree in each step is often not computationally tractable for multivariate data. Instead one can adopt a greedy, coarse-to-fine algorithm in which we iteratively choose a single cut point at each split of the tree $T_{k}$ that achieves the steepest ascent in the Hellinger distance or the steepest ascent in the corresponding affinity. Specifically, when we consider splitting the current node $A$, we can search for the splitting rule such that for the child nodes $A_{l}$ and $A_{r}$ the following affinity is minimized:
@@
$\frac{\partial l_{n}(w)}{\partial F(x_{i}^{(0)})}=-\frac{1}{n_{0}}e^{-F(x_{i}^{(0)})}=\frac{1}{n_{0}}w^{-1}(x_{i}^{(0)})\quad\text{and}\quad-\frac{\partial l_{n}(w)}{\partial F(x_{i}^{(1)})}=\frac{1}{n_{1}}e^{F(x_{i}^{(1)})}=-\frac{1}{n_{1}}w(x_{i}^{(1)}),$
which are the pseudo-residuals. One can then follow the recipe provided in (Friedman, 2001). Specifically, in each iteration of the algorithm we fit a regression tree to the pseudo-residuals by minimizing the squared error loss, that is, minimizing the variances of the gradients in the leaf nodes. We again adopt a coarse-to-fine greedy search and for each node $A$, we select the splitting rule with child nodes $A_{l}$ and $A_{r}$ that minimizes the variances. After constructing the tree, we obtain the optimal values of the node parameters, which are already given by Proposition 2.1.
@@
An additional, complementary strategy for regularization is to introduce a restriction on the size of trees to avoid constructing overly complicated partition structures, so we set the maximum resolution (depth) of the trees to be, for example, between 4 and 6. Finally, the number of trees in the ensemble also needs to be carefully selected. We find that choosing this number via cross-validation yields satisfactory performance.
@@ 3.1 Balancing loss as a pseudo-likelihood
Tree boosting only offers point estimates for the density ratio function, without providing any uncertainty associated. To achieve uncertainty quantification for the density ratio, we can instead adopt a (generalized) Bayesian strategy (Bissiri et al., 2016) to learning the density ratio. Specifically, we define a pseudo-likelihood using the finite-sample loss $l_{n}(w)$, that is
$L_{n,\tau}(w)=\exp(-n\tau l_{n}(w)),$
where $\tau>0$ is a “temperature” parameter. Then, given the prior distribution of $w$, $\pi_{0}(w)$, the posterior distribution of $w$ is defined as
@@ 3.2 Conjugate prior specification
Next we describe prior specification on $w$ for efficient MCMC and a data-driven approach for setting the value of $\tau$. Recalling our model introduced in (4), we describe a Gibbs sampling strategy that iteratively samples the full conditional of each basis function $f_{k}$, given all the others. That is, how one may sample the node parameters $\beta_{k}(A)$ and the tree partition $T_{k}$ that define $f_{k}$ as in (5).
3.2.1 Updating leaf node parameters
@@
$L_{n,\tau}(w)\propto\exp\left(-\zeta^{-1}\tau\gamma_{k}(A)^{-1}\sum_{x^{(0)}_{i}\in A}w_{-k}^{-1}(x^{(0)}_{i})-(1-\zeta)^{-1}\tau\gamma_{k}(A)\sum_{x^{(1)}_{i}\in A}w_{-k}(x^{(1)}_{i})\right)$,
where $\zeta=n_{0}/n$ and $\gamma_{k}(A)=e^{\beta_{k}(A)}$. We note that this pseudo-likelihood resembles an exponential family kernel for $\gamma_{k}(A)$, and as such we adopt an inverse-Gaussian prior for $\gamma_{k}(A)$, which is conjugate to this kernel. Specifically, the prior $\mathrm{InverseGauss}(\mu_{k}(A),\lambda_{k}(A))$ on $\gamma_{k}(A)$ takes the following form:
@@
$\mu^{\prime}_{k}(A)=\sqrt{\frac{\lambda_{k}(A)+2\zeta^{-1}\tau\sum_{x^{(0)}_{i}\in A}w_{-k}^{-1}(x^{(0)}_{i})}{\frac{\lambda_{k}(A)}{\mu_{k}(A)^{2}}+2(1-\zeta)^{-1}\tau\sum_{x^{(1)}_{i}\in A}w_{-k}(x^{(1)}_{i})}},\ \lambda^{\prime}_{k}(A)=\lambda_{k}(A)+2\zeta^{-1}\tau\sum_{x^{(0)}_{i}\in A}w_{-k}^{-1}(x^{(0)}_{i})$. (6)
We set $\mu_{k}(A)$ to $1$ so that the induced marginal prior on $\log w$ is around 0 (more discussions are to be given at the end of this section). The precision parameter $\lambda_{k}(A)$ induces shrinkage toward the prior mean 1, as seen in the posterior mean $\mu^{\prime}_{k}(A)$ given in Eq (6). We set $\lambda_{k}(A)$ according to the number of trees $K$:
$\lambda_{k}(A)=\lambda_{0}K,$
where $\lambda_{0}$ is a hyperparameter. The prior value of $\lambda_{k}(A)$ is linear in $K$ because the results in Whitmore and Yalovsky (1978) imply that the variance of $f_{k}(x)$ for each point $x$ ($x\in\Omega$) is $O(K^{-1})$ under our model specification (see Supplementary Materials A for more details). This choice of $\lambda_{k}(A)$ is also analogous to the amount of shrinkage recommended for the BART model (Chipman et al., 2010). We set $\lambda_{0}$ to 5 as a default value in all of our numerical experiments.
Another consideration is that the inverse Gaussian distribution has an asymmetric shape, so it can also introduce skewness into the prior of the log-balancing weight $\log w$. To avoid this issue, we apply the inverse Gaussian prior to $\gamma_{k}(A)$ and $\gamma^{-1}_{k,A}$ alternately, that is,
$\gamma_{k}(A)$ $\sim\mathrm{InverseGauss}(1,\lambda_{0}K)\ (k\text{: odd}),$
$\gamma_{k}(A)^{-1}$ $\sim\mathrm{InverseGauss}(1,\lambda_{0}K)\ (k\text{: even}).$
@@ 3.3 Tuning the temperature $\tau$
In generalized Bayesian inference, the temperature parameter $\tau$ controls the strength of the loss-based likelihood for the balancing weight function $w$ relative to the prior and significantly influences the resulting posteriors, including posterior variances and credible intervals. There is a growing body of literature on strategies for selecting the temperature parameter empirically. Some proposed methods include information matching (Holmes and Walker, 2017), calibrating frequentist coverage rates of credible intervals (Syring and Martin, 2019), and the SafeBayes algorithm (Grünwald and van Ommen, 2017). However, it is not straightforward to adapt these methods to the current context, where the task is unsupervised and the function of interest is an unknown, nonparametrically specified function.
@@
Instead, we adopt a hierarchical Bayesian approach by placing a prior on the temperature $\tau$, following Bissiri et al. (2016); Rigon et al. (2023). To this end, we add $\tau^{n}$ to the pseudo-likelihood so that the full conditional of $\tau$ is proportional to
$\pi_{0}(\tau)\tau^{n}\prod^{n_{0}}_{i=1}\exp\left(-\zeta^{-1}\tau w^{-1}(x^{(0)}_{i})\right)\prod^{n_{1}}_{i=1}\exp\left(-(1-\zeta)^{-1}\tau w(x^{(1)}_{i})\right)$,
where $\pi_{0}(\tau)$ denotes the prior for $\tau$. We adopt a conjugate prior $\pi_{0}(\tau)$, namely $\mathrm{Gamma}(a_{0,\tau},b_{0,\tau})$, so the full conditional of $\tau$ is $\mathrm{Gamma}(a_{1,\tau},b_{1,\tau})$, where
$a_{1,\tau}=a_{0,\tau}+n,\ b_{1,\tau}=b_{0,\tau}+nl_{n}(w).$
The analytic form of the full conditional elucidates the role $\tau$ plays and justifies the data-dependent approach to tuning its value. Specifically, $\tau$ can be interpreted as a regularization parameter that penalizes poor overlap of the two samples. To see this, note that the full conditional mean of $\tau$ is $\frac{a_{0,\tau}+n}{b_{0,\tau}+nl_{n}(w)}$, so when the total sample size is large, the inverse of this quantity approximates the realized value of the loss $l_{n}(w)$. When the two samples have little overlap, for example, the estimated weights $w^{-1}$ for the first group and $w$ for the second group tend to be both small and close to zero, leading to a small $l_{n}(w)$ and thus a large $\tau$. Conversely, when the samples have good overlap, the weights are closer to 1, yielding a larger $l_{n}(w)$ and a smaller $\tau$. As an illustration, Figure 1 shows the posterior distributions of $\tau$ under different levels of overlap under a simple simulation experiment. The observations ($n_{0}=n_{1}=200$) are generated from the standard normal distribution, and the second group is mean-shifted by $1,2$, and $3$, corresponding to decreasing levels of overlap. We can see that poorer overlap of the two samples leads to larger absolute values of $\log w$, along with larger values of $\tau$ corresponding to the smaller realized value of the loss $l_{n}(w)$.
Figure 1: Posterior distributions of $\tau$ under different levels of overlap. Top row: The black and red points correspond to the two samples, and the dashed lines indicate the posterior means of the log-density ratios. Bottom row: The density functions of the posterior distributions of $\tau$ are shown.

### v3

@@ 2.3 Additive Tree Models and Boosting
Given the close relationship between density ratios and posterior odds, the latter of which can be estimated using supervised tree boosting with additive tree models, it is natural to consider the class of additive tree models as a promising candidate for approximating the density ratio. We thus consider from now on fitting additive tree models to the density ratio under the balancing loss. In this section, we begin with a non-Bayesian approach that aims to find a (functional) point estimate of the density ratio by optimizing the finite-sample version of the balancing loss. The resulting training algorithms are very similar to those used in supervised tree boosting.
Before introducing the algorithms, we describe the additive tree model. We consider a model for the log-balancing weight $\log w$ in the form
@@
Proposition 1.
Suppose $w_{k-1}$ is the estimate of $\sqrt{r^{*}}=\sqrt{p/q}$ after (k1)(k-1) steps of fitting. Then in the $k$th step, the loss is minimized by $\log w_{k}=\log w_{k-1}+f_{k}$ when the following conditions are satisfied by $f_{k}$:
@@
    $\exp(\beta_{k}(A))=\sqrt{\frac{\int_{A}pw^{-1}_{k-1}\,d\mu}{\int_{A}qw_{k-1}\,d\mu}}.$
Remark 1: If the balancing weight is correctly estimated, it achieves the symmetry between the two expectations in the loss in that $\mathbb{E}_{p}[w^{-1}]=\mathbb{E}_{q}[w]$, and we can impose this condition as a restriction or regularization on the estimator so that it does not focus on minimizing only one expectation out of the two. As justification, Proposition 1 implies that the balancing loss can be improved by multiplying the current estimate $w_{k}$ by a constant ($c$, say) such that $\mathbb{E}_{p}[(cw_{k})^{-1}]=\mathbb{E}_{q}[cw_{k}]$ (see Supplementary Materials A for details). Based on this discussion, we introduce this correction step in the boosting algorithms.
@@
As in tree-based predictive models, searching for the exact optimal tree in each step is often not computationally tractable for multivariate data. Instead one can adopt a greedy, coarse-to-fine algorithm in which we iteratively choose a single cut point at each split of the tree $T_{k}$ that achieves the steepest ascent in the Hellinger distance or the steepest ascent in the corresponding affinity. Specifically, when we consider splitting the current node $A$, we can search for the splitting rule such that for the child nodes $A_{l}$ and $A_{r}$ the following Hellinger affinity is minimized, or, equivalently, the Hellinger distance is maximized:
@@
$\frac{\partial l_{n}(w)}{\partial F(x_{i}^{(0)})}=-\frac{1}{n_{0}}\exp(-F(x_{i}^{(0)}))=\frac{1}{n_{0}}w^{-1}(x_{i}^{(0)}),\ -\frac{\partial l_{n}(w)}{\partial F(x_{i}^{(1)})}=\frac{1}{n_{1}}\exp(F(x_{i}^{(1)}))=-\frac{1}{n_{1}}w(x_{i}^{(1)}),$
which are the pseudo-residuals. One can then follow the recipe provided in (Friedman, 2001). Specifically, in each iteration of the algorithm we fit a regression tree to the pseudo-residuals by minimizing the squared error loss, that is, minimizing the variances of the gradients in the leaf nodes. We again adopt a coarse-to-fine greedy search and for each node $A$, we select the splitting rule with child nodes $A_{l}$ and $A_{r}$ that minimizes the variances. After constructing the tree, we obtain the optimal values of the node parameters, which are already given by Proposition 1.
@@
An additional, complementary strategy for regularization is to introduce a restriction on the size of trees to avoid constructing overly complicated partition structures, so we set the maximum resolution (depth) of the trees to be, for example, 4. Finally, the number of trees in the ensemble also needs to be carefully selected. We find that choosing this number via cross-validation yields satisfactory performance.
@@ 3 Generalized Bayesian Inference on Additive Trees for Density Ratios
Tree boosting only offers point estimates for the density ratio function, without providing any associated uncertainty. To achieve uncertainty quantification for the density ratio, in this section, we introduce a generalized Bayesian strategy (Bissiri et al., 2016) to learning the density ratio.
@@ 3.1 Balancing Loss as a Pseudo-likelihood
Following the framework introduced in (Bissiri et al., 2016), we define a pseudo-likelihood using the finite-sample loss $l_{n}(w)$, that is,
$L_{n,\tau}(w)=\frac{\exp(-n_{\mathrm{min}}\tau l_{n}(w))}{\tau^{2n_{\mathrm{min}}}}$
where $n_{\min}$ is the sample size of the smaller group $\min(n_{0},n_{1})$, and $\tau>0$ is a “temperature” parameter. The term in the denominator $\tau^{2n_{\mathrm{min}}}$ is added to provide a clearer interpretation of the pseudo-likelihood as a weighted likelihood under exponential distributions, as follows:
$L_{n,\tau}(w)=\prod^{n_{0}}_{i=1}\left[\tau\exp\left(-\tau w^{-1}(x^{(0)}_{i})\right)\right]^{\zeta_{0}}\times\prod^{n_{1}}_{i=1}\left[\tau\exp\left(-\tau w(x^{(1)}_{i})\right)\right]^{\zeta_{1}},$ (4)
where $\zeta_{0}=n_{\mathrm{min}}/n_{0}$ and $\zeta_{1}=n_{\mathrm{min}}/n_{1}$ are the weights taking values in $(0,1]$. If, for example, the first group has a smaller sample size with $n_{0}<n_{1}$, then the observations from the second group are down-weighted to keep the information from the two groups balanced. The effective sample size is controlled not by the total sample size but by $n_{\mathrm{min}}$, and this is natural since the amount of information about the two-sample difference brought by the data crucially depends on the size of the smaller group.
Under the formulation, the temperature $\tau$ can also be interpreted by a one-dimensional summary of the two-sample difference. To see this, we note that the likelihood $L_{n,\tau}(w)$ is maximized with respect to $\tau$ when $\tau^{-1}=\frac{1}{2}l_{n}(w)$. As $w$ is the balancing function corresponding to $\sqrt{p/q}$, what we estimate with $\tau^{-1}$ is
$\frac{1}{2}l_{n}(\sqrt{p/q})=\int\sqrt{pq}\,d\mu,$
which is the Bhattacharyya coefficient (Bhattacharyya, 1946).
Given the prior distribution of $w$, $\pi_{0}(w)$, the posterior distribution of $w$ is defined as
@@
where $\pi_{0}(w)$ is the prior of the balancing function $w$. Its specification is detailed in the next section.
@@ 3.2 Conjugate Prior Specification
Next we describe prior specification on $w$ for efficient MCMC and a data-driven approach for setting the value of $\tau$. Recalling our model introduced in Eq. (2), we describe a Gibbs sampling strategy that iteratively samples from the full conditional of each basis function $f_{k}$, given all the others. That is, how one may update the node parameters $\beta_{k}(A)$ and the tree partition $T_{k}$ that define $f_{k}$ defined in Eq. (3).
3.2.1 Updating The Leaf Node Parameters
@@
$L_{n,\tau}(w)\propto\exp\left(-{\color[rgb]{0,0,0}{\zeta_{0}}}\tau\gamma_{k}(A)^{-1}\sum_{x^{(0)}_{i}\in A}w_{-k}^{-1}(x^{(0)}_{i})-{\color[rgb]{0,0,0}{\zeta_{1}}}\tau\gamma_{k}(A)\sum_{x^{(1)}_{i}\in A}w_{-k}(x^{(1)}_{i})\right)$,
where $\gamma_{k}(A)=\exp(\beta_{k}(A))$. Noting that this pseudo-likelihood resembles an exponential family kernel for $\gamma_{k}(A)$, we adopt an inverse-Gaussian leaf prior (Murray, 2021) for $\gamma_{k}(A)$, which is conjugate to this kernel. Specifically, the prior $\mathrm{InverseGauss}(\mu_{k}(A),\lambda_{k}(A))$ on $\gamma_{k}(A)$ takes the following form:
@@
$\mu^{\prime}_{k}(A)=\sqrt{\frac{\lambda_{k}(A)+2{\color[rgb]{0,0,0}{\zeta_{0}}}\tau\sum_{x^{(0)}_{i}\in A}w_{-k}^{-1}(x^{(0)}_{i})}{\frac{\lambda_{k}(A)}{\mu_{k}(A)^{2}}+2{\color[rgb]{0,0,0}{\zeta_{1}}}\tau\sum_{x^{(1)}_{i}\in A}w_{-k}(x^{(1)}_{i})}},\ \lambda^{\prime}_{k}(A)=\lambda_{k}(A)+2{\color[rgb]{0,0,0}{\zeta_{0}}}\tau\sum_{x^{(0)}_{i}\in A}w_{-k}^{-1}(x^{(0)}_{i})$.
We set $\mu_{k}(A)$ to $1$ so that the induced marginal prior on $\log w$ is around 0 (more discussions are to be given at the end of this section). The precision parameter $\lambda_{k}(A)$ induces shrinkage toward the prior mean 1, as seen in the expression of the posterior mean $\mu^{\prime}_{k}(A)$. We set $\lambda_{k}(A)$ according to the number of trees $K$, namely, $\lambda_{k}(A)=\lambda_{0}K,$ where $\lambda_{0}$ is a hyperparameter. The prior value of $\lambda_{k}(A)$ is linear in $K$ because the results in Whitmore and Yalovsky (1978) imply that the variance of $f_{k}(x)$ for each point $x$ ($x\in\Omega$) is $O(K^{-1})$ under our model specification (see Supplementary Materials A for more details). This choice of $\lambda_{k}(A)$ is also analogous to the amount of shrinkage recommended for the BART model (Chipman et al., 2010). We set $\lambda_{0}$ to 5 as a default value in all of our numerical experiments.
Another consideration is that the inverse Gaussian distribution has an asymmetric shape, so it can also introduce skewness into the prior of the log-balancing weight $\log w$. To avoid this issue, we apply the inverse Gaussian prior to $\gamma_{k}(A)$ and $\gamma_{k}(A)^{-1}$ alternately, that is,
$\gamma_{k}(A)\sim\mathrm{InverseGauss}(1,\lambda_{0}K)\ (k\text{: odd}),\ \gamma_{k}(A)^{-1}\sim\mathrm{InverseGauss}(1,\lambda_{0}K)\ (k\text{: even}).$
@@ 3.3 Tuning the Temperature $\tau$
In generalized Bayesian inference, the temperature parameter $\tau$ controls the strength of the loss-based likelihood for the balancing weight function $w$ relative to the prior and significantly influences the resulting posteriors, including posterior variances and credible intervals. There is a growing body of literature on strategies for selecting the temperature parameter empirically. Some proposed methods include information matching (Holmes and Walker, 2017), calibrating frequentist coverage rates of credible intervals (Syring and Martin, 2019), and the SafeBayes algorithm (Grünwald and van Ommen, 2017). However, it is not straightforward to adapt these methods to the current context, where the task is to learn of an unknown, nonparametrically specified function in an unsupervised setting and the computation cost of the MCMC is relatively expensive.
@@
Instead, we adopt a hierarchical Bayesian approach by placing a prior on the temperature $\tau$, following Bissiri et al. (2016); Rigon et al. (2023). Motivated by the weighted likelihood expression in Eq. (4), we adopt a conjugate prior $\pi_{0}(\tau)$, namely $\mathrm{Gamma}(a_{0,\tau},b_{0,\tau})$, so the full conditional of $\tau$ is $\mathrm{Gamma}(a_{1,\tau},b_{1,\tau})$, where $a_{1,\tau}=a_{0,\tau}+2n_{\mathrm{min}},\ b_{1,\tau}=b_{0,\tau}+n_{\mathrm{min}}l_{n}(w).$
3.4 Empirical Validation of Uncertainty Calibration
We provide an illustration of the generalized Bayesian inference under two one-dimensional Gaussian distributions, $N(0,1)$ and $N(1,2.25)$, to verify whether the resulting generalized Bayesian uncertainty is properly calibrated.
The posterior distributions under the balanced and unbalanced sample sizes are provided in Figure 1, where the posterior means and the credible intervals are evaluated pointwise. We can see that the posterior distributions are distributed around the true log-densities in a region with an adequate number of observations (i.e., the region between -2 and 2). Moreover, the inverse temperature $\tau^{-1}$, with which we estimate the Bhattacharyya coefficient, is distributed around the true value. We note that more information about the two-sample difference is provided under the balanced sample size as is represented by the weighted likelihood expression, and it is reflected by the different widths of the credible intervals in the posterior distributions. Figure 1 also shows the calibration of the credible intervals obtained by our Bayesian model in terms of nominal coverage rates based on 50 simulated data sets. (For each simulation, the coverage rate is calculated based on the ratio of the observations included in the corresponding credible intervals.) It implies that the posterior uncertainty and the nominal coverage rate tend to agree more under a larger sample size, as in the case of $(n_{0},n_{1})=(2500,2500)$. It is noteworthy that the results under $(500,500)$ and $(500,4500)$ are similar, and again it is consistent with the specification of the loss-based likelihood, in which the amount of information is controlled by the size of the smaller group ($n_{0}=500$ in this example).
(a) Balanced ($n_{0}=500,n_{1}=500$)
(b) Unbalanced ($n_{0}=100,n_{1}=900$)
(c) Balanced ($n_{0}=2500,n_{1}=2500$)
(d) Unbalanced ($n_{0}=500,n_{1}=4500$)
Figure 1: An evaluation of posterior distributions under the 1D scenarios with the balanced/unbalanced sample size. Each left plot compares the true log-densities (gray, dotted) and the posteriors (mean: black, solid, 95% interval: blue, dashed) evaluated pointwise. Each middle plot shows the posterior distributions of the inverse temperature $\tau^{-1}$ and the true Bhattacharyya coefficient. Each right plot visualizes the calibration plot comparing the posterior uncertainty and the nominal coverage rate (namely, the ratio of observations included in the pointwise credible intervals) based on 50 simulations and their averages.